# Game Theory in Cosmology

### v1

@@ 4.5 Teleodynamic Cosmological Entropy
$p_{i}=\frac{1}{z_{\text{cell}}}\exp[-\lambda_{\tiny E}E_{i}-\lambda_{\tiny\Phi}\Phi_{i}]$ (75)
@@
$Z(H)=[z_{\text{cell}}(\lambda_{E},\lambda_{\Phi})]^{N(H)}$ (76)
@@
$S_{\text{\tiny TD}}(H)=-\sum_{\{i\}}p_{i}\ln p_{i}=N(H)s_{\text{cell}}(\lambda_{\tiny E},\lambda_{\tiny\Phi})$ (77)
@@
$S_{\text{\tiny TD}}(H)=\frac{4\pi s_{\text{cell}}}{\ell_{\chi}^{2}}\frac{1}{H^{2}}\equiv\frac{C_{\text{\tiny TD}}}{H^{2}}$ (78)
@@
$C_{\text{\tiny TD}}=\frac{4\pi s_{\text{cell}}}{\ell_{\chi}^{2}}$ (79)
@@
$\sigma_{\text{\tiny TD}}(t)=\sum_{a}L_{ab}(t)J_{a}(t)J_{b}(t)\geq 0$ (80)
@@
$S_{\text{\tiny TD}}(t)=\frac{C_{\text{\tiny TD}}}{H(t)^{2}}+\int^{t}\sigma_{\text{\tiny TD}}(t^{\prime})dt^{\prime}$ (81)
@@
$T^{-1}(t)=\frac{\partial S_{\text{\tiny TD}}}{\partial E_{H}}$ (82)
@@
$E_{H}(t)=\frac{4\pi}{3}r_{H}^{3}\rho_{\text{eff}}=\frac{4\pi M_{P}^{2}}{H(t)}$ (83)
@@
$\frac{dE_{H}}{dH}=-\frac{4\pi M_{P}^{2}}{H^{2}},\quad\frac{dS_{\text{\tiny TD}}}{dH}=-\frac{2C_{\text{\tiny TD}}}{H^{3}}$ (84)
@@
$T_{\text{\tiny TD}}(H)=\left(\frac{\partial S_{\text{\tiny TD}}}{\partial E_{H}}\right)^{-1}=\frac{2\pi M_{P}^{2}}{C_{\text{\tiny TD}}}H$ (85)
@@ 4.6 Observational Signatures of Cosmological Teleodynamics
We next consider the observational consequences of teleodynamics. Since this framework is based on nonlocal memory and environmental correlations rather than additional particle fields, several of its signatures cannot be reproduced by conventional dark sector models, which is an advantage. These include distinctive patterns in galactic halo profiles, anomalous behavior of peculiar velocities and scale-dependent modifications in large scale structure growth.
@@
$\varphi(x)=T^{ij}(x)\frac{\partial_{i}\partial_{j}}{\nabla^{2}}\delta(x)$ (87)
@@
$\mathbf{F}_{\text{\tiny TD}}(x)=-\alpha\nabla\varphi(x)=-\alpha T^{ij}\nabla_{i}\left[\frac{\partial_{j}}{\nabla^{2}}\delta(x)\right]$ (88)
@@
$\Psi_{\text{eff}}(x)=\Psi(x)+\alpha T^{ij}\frac{\partial_{i}\partial_{j}}{\nabla^{2}}\delta(x)$ (89)
@@
$\nabla^{2}\Psi_{\text{eff}}\propto\delta+\alpha T^{ij}\hat{k}_{i}\hat{k}_{j}\delta(k)$ (90)
@@
$\theta(k,a)=-aH[1+\Delta\mu(k,a)]\delta(k,a)$ (93)
@@
$u_{i}(x,a)=-aH\int\frac{d^{3}k}{(2\pi)^{3}}e^{ik\cdot x}\left[\hat{k}_{i}+\Delta\mu(k,a)T_{ij}(\hat{k})\hat{k}_{j}\right]\frac{\delta(k,a)}{k}$ (94)
@@
$\langle u^{2}(R)\rangle=a^{2}H^{2}\int\frac{d^{3}k}{(2\pi)^{3}}[1+\Delta\mu(k,a)]^{2}\frac{P_{\delta\delta}(k,a)}{k^{2}}W^{2}(kR)$ (95)
@@
$\rho_{\text{TD}}(k,a)=\Delta\mu(k,a)\bar{\rho}\delta(k,a)$ (96)
@@ 5 Law of Universal Arbitrage Equilibrium
$S_{\rm TD}(t)=\frac{C_{\rm TD}(t)}{H(t)^{2}}+\int^{t}\sigma_{\rm TD}(t^{\prime})\,dt^{\prime},\qquad T_{\rm TD}(t)=\left(\frac{\partial S}{\partial E_{H}}\right)^{-1}\simeq\frac{2\pi M_{P}^{2}}{C_{\rm TD}(t)}H(t)$ (97)
@@
$\dot{S}_{\rm TD}=-\frac{2C_{\rm TD}\dot{H}}{H^{3}}+\frac{\dot{C}_{\rm TD}}{H^{2}}+\sigma_{\rm TD}$ (98)
@@
$\dot{H}=-\frac{\rho_{\rm eff}+p_{\rm eff}}{2M_{P}^{2}}+\frac{H}{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+\frac{H^{3}}{2C_{\rm TD}}\,\sigma_{\rm TD}$ (101)
@@
Now, if we go by multiplying the equation (101) by $2H$ and using $\tfrac{d}{dt}(H^{2})=2H\dot{H}$ together with the continuity equation $\dot{\rho}_{\rm eff}+3H(\rho_{\rm eff}+p_{\rm eff})=0$ then this leads us to
$\frac{d}{dt}(H^{2})=\frac{1}{3M_{P}^{2}}\dot{\rho}_{\rm eff}+H^{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+H^{4}\frac{\sigma_{\rm TD}}{C_{\rm TD}}$ (102)
@@
$H^{2}(t)=\frac{\rho_{\rm eff}(t)}{3M_{P}^{2}}+\left[H^{2}(t_{\star})-\frac{\rho_{\rm eff}(t_{\star})}{3M_{P}^{2}}\right]+\int_{t_{\star}}^{t}\left(H^{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+H^{4}\frac{\sigma_{\rm TD}}{C_{\rm TD}}\right)dt^{\prime}$ (103)
@@
In this language, equation (101) becomes
$\dot{H}=-\frac{\rho_{m}+\rho_{r}+p_{r}}{2M_{P}^{2}}-\frac{\rho_{\rm TD}+p_{\rm TD}}{2M_{P}^{2}}+\frac{H}{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+\frac{H^{3}}{2C_{\rm TD}}\,\sigma_{\rm TD}$ (104)
revealing that two distinct teleodynamic mechanisms contribute to the expansion rate with a bulk component carried by $(\rho_{\rm TD},p_{\rm TD})=\alpha(\overline{\Phi},\overline{\Pi}_{\rm TD})$ and a horizon component encoded in $C_{\rm TD}(t)$ and $\sigma_{\rm TD}(t)$ representing the evolution of the correlation scale and the non-equilibrium memory production on the horizon. If late-time acceleration is primarily driven by bulk bias, then $C_{\rm TD}$ remains approximately constant and $\sigma_{\rm TD}$ approaches zero, with $w_{\rm TD}\approx-1$ emerging naturally when $\overline{\Pi}_{\rm TD}\simeq-\overline{\Phi}$. If instead part of the acceleration originates from the horizon sector, then the integral terms in the teleodynamic Friedmann relation are nonzero, and one infers small but finite values for $\dot{C}_{\rm TD}/C_{\rm TD}$ or $\sigma_{\rm TD}$ at low redshift. Since $\sigma_{\rm TD}\geq 0$, the last term in (101) always drives $H^{2}$ upward, temporarily mimicking an effective equation of state below 1-1 without introducing ghosts.
@@
$\dot{S}_{\rm TD}=-2C_{\rm TD}\frac{\dot{H}}{H^{3}}+\sigma_{\rm TD},\qquad\dot{E}_{H}=-4\pi M_{P}^{2}\frac{\dot{H}}{H^{2}},\qquad\dot{V}_{H}=-4\pi\frac{\dot{H}}{H^{4}}$ (106)
@@
$p_{\rm eff}=\frac{\kappa}{4\pi}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD},\qquad\kappa=\frac{2\pi M_{P}^{2}}{C_{\rm TD}}$ (107)
@@
$\alpha\overline{\Pi}_{\rm TD}=\frac{\kappa}{4\pi}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}=\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}$ (108)
@@
$\alpha\overline{\Phi}=-2M_{P}^{2}\dot{H}-(\rho_{m}+p_{m}+\rho_{r}+p_{r})-\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}$ (109)
@@
$p_{\rm TD}=\alpha\overline{\Pi}_{\rm TD}=-\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{|\dot{H}|}\sigma_{\rm TD}<0$
showing that non-equilibrium teleodynamic entropy production necessarily gives us negative pressure and a combined expression for the EOS parameter follows
$w_{\rm TD}=\frac{p_{\rm TD}}{\rho_{\rm TD}}=\frac{-\dfrac{M_{P}^{2}}{2C_{\rm TD}}\dfrac{H^{5}}{\dot{H}}\sigma_{\rm TD}}{-2M_{P}^{2}\dot{H}-(\rho_{m}+p_{m}+\rho_{r}+p_{r})-\dfrac{M_{P}^{2}}{2C_{\rm TD}}\dfrac{H^{5}}{\dot{H}}\sigma_{\rm TD}}$ (110)
@@
$\sigma_{\rm TD}(t)=2\xi\frac{-\dot{H}}{H^{3}},\qquad\xi>0$ (111)
@@
$p_{\rm TD}=-\frac{\kappa}{2\pi}\xi H^{2}=-\frac{M_{P}^{2}}{C_{\rm TD}}\xi H^{2},\qquad\rho_{\rm TD}=\frac{3M_{P}^{2}H^{2}}{8\pi}\Big(\cdots\Big)+\frac{M_{P}^{2}}{C_{\rm TD}}\xi H^{2}$ (112)
@@
depends exclusively on horizon-level quantities such as $C_{\rm TD}$, $\sigma_{\rm TD}$, $H$, and $\dot{H}$ and none of the spatially varying components of the teleodynamic memory field enter this expression, indicating that cosmic acceleration arises only from the homogeneous part of the memory sector. Global memory, therefore, generates dark energy behavior while local memory enters through spatial gradients such as $\alpha\nabla\varphi$ and drives clustering, thereby reproducing dark matter like dynamics.
The form of the pressure above also shows that because $\sigma_{\rm TD}\geq 0$ and $\dot{H}<0$ in an expanding Universe, the negative pressure driving acceleration emerges if and only if the accessible microstates at the horizon increase. Teleodynamics, therefore, provides a thermodynamic demonstration and a law.

### v2

@@ 5.5 Teleodynamic Cosmological Entropy
$p_{i}=\frac{1}{z_{\text{cell}}}\exp[-\lambda_{\tiny E}E_{i}-\lambda_{\tiny\Phi}\Phi_{i}]$ (5.65)
@@
$Z(H)=[z_{\text{cell}}(\lambda_{E},\lambda_{\Phi})]^{N(H)}$ (5.66)
@@
$S_{\text{\tiny TD}}(H)=-\sum_{\{i\}}p_{i}\ln p_{i}=N(H)s_{\text{cell}}(\lambda_{\tiny E},\lambda_{\tiny\Phi})$ (5.67)
@@
$S_{\text{\tiny TD}}(H)=\frac{4\pi s_{\text{cell}}}{\ell_{\chi}^{2}}\frac{1}{H^{2}}\equiv\frac{C_{\text{\tiny TD}}}{H^{2}}$ (5.68)
@@
$C_{\text{\tiny TD}}=\frac{4\pi s_{\text{cell}}}{\ell_{\chi}^{2}}$ (5.69)
@@
$\sigma_{\text{\tiny TD}}(t)=\sum_{a}L_{ab}(t)J_{a}(t)J_{b}(t)\geq 0$ (5.70)
@@
$S_{\text{\tiny TD}}(t)=\frac{C_{\text{\tiny TD}}}{H(t)^{2}}+\int^{t}\sigma_{\text{\tiny TD}}(t^{\prime})dt^{\prime}$ (5.71)
@@
$T^{-1}(t)=\frac{\partial S_{\text{\tiny TD}}}{\partial E_{H}}$ (5.72)
@@
$E_{H}(t)=\frac{4\pi}{3}r_{H}^{3}\rho_{\text{eff}}=\frac{4\pi M_{P}^{2}}{H(t)}$ (5.73)
@@
$\frac{dE_{H}}{dH}=-\frac{4\pi M_{P}^{2}}{H^{2}},\quad\frac{dS_{\text{\tiny TD}}}{dH}=-\frac{2C_{\text{\tiny TD}}}{H^{3}}$ (5.74)
@@
$T_{\text{\tiny TD}}(H)=\left(\frac{\partial S_{\text{\tiny TD}}}{\partial E_{H}}\right)^{-1}=\frac{2\pi M_{P}^{2}}{C_{\text{\tiny TD}}}H$ (5.75)
@@
We would like to highlight a subtle point in passing here. The "microstate counting" that appears in the teleodynamic derivation of cosmological horizon entropy is conceptually distinct from the microscopic state counting performed in string-theoretic black hole entropy [73, 74, 75]. In teleodynamics, the quantity $N_{\rm TD}(t)=A_{\rm hor}(t)/\ell_{\chi}^{2}(t)$ counts coarse-grained correlation cells of size $\ell_{\chi}(t)$ on a time-dependent apparent horizon and the entropy $S_{\rm TD}(t)=N_{\rm TD}(t)\,s_{\rm cell}(t)+\int^{t}dt^{\prime}\,\sigma_{\rm TD}(t^{\prime})$ reflects both the instantaneous coarse-grained structure and the accumulated non-equilibrium memory encoded in $\sigma_{\rm TD}(t)$. Note that the procedure here is primarily semi-classical, and these cells are not quantum microstates in a Hilbert space, as they are classical correlation domains whose number and entropy evolve with $H(t)$, $\ell_{\chi}(t)$, and structure formation. By contrast, string-theoretic microstate counting for black holes counts exact quantum states in a stationary sector [76, 77, 78, 79, 80, 81] labeled by fixed charges $(M,J,Q)$, whose degeneracy is protected by supersymmetry or modular invariance. This counting requires a Killing horizon, a time-translation generator, and a static Hilbert-space partition, none of which exist for a cosmological horizon.
This brings us to another central point of our work here. If a complete quantum gravity microstate formula for cosmology exists, its semiclassical limit therefore cannot reproduce the stationary Bekenstein–Hawking area law. Instead, it must reduce to the teleodynamic entropy $S_{\rm TD}(t)$, because cosmology is intrinsically non-stationary, with $\dot{H}\neq 0$, evolving correlation lengths and non-zero entropy production $\sigma_{\rm TD}(t)$. The universe does not admit a fixed charge and a time-independent microcanonical sector, so the semiclassical horizon entropy cannot be invariant under $H\to H(t)$. Any correct QG construction must yield $S_{\rm QG}^{\rm semiclassical}(t)\to S_{\rm TD}(t)$ and not $A/4G$, which would end up reflecting that cosmological horizons live in the non-equilibrium, memory-bearing sector of gravitational thermodynamics, while black holes inhabit the equilibrium limit in which teleodynamic corrections collapse to constants. This provides a semi-classical interpretation of the thermodynamic split conjecture [82, 83] and highlights teleodynamics as the appropriate semiclassical bridge between cosmology and quantum gravity.
@@ 5.6 Observational Signatures of Cosmological Teleodynamics
We next consider the observational consequences of teleodynamics. Since this framework is based on nonlocal memory and environmental correlations rather than additional particle fields, several of its signatures cannot be reproduced by conventional dark sector models, which is an advantage. These include distinctive patterns in galactic halo profiles, anomalous behavior of peculiar velocities, and scale-dependent modifications in large-scale structure growth.
@@
$\varphi(x)=T^{ij}(x)\frac{\partial_{i}\partial_{j}}{\nabla^{2}}\delta(x)$ (5.77)
@@
$\mathbf{F}_{\text{\tiny TD}}(x)=-\alpha\nabla\varphi(x)=-\alpha T^{ij}\nabla_{i}\left[\frac{\partial_{j}}{\nabla^{2}}\delta(x)\right]$ (5.78)
@@
$\Psi_{\text{eff}}(x)=\Psi(x)+\alpha T^{ij}\frac{\partial_{i}\partial_{j}}{\nabla^{2}}\delta(x)$ (5.79)
@@
$\nabla^{2}\Psi_{\text{eff}}\propto\delta+\alpha T^{ij}\hat{k}_{i}\hat{k}_{j}\delta(k)$ (5.80)
@@
$\theta(k,a)=-aH[1+\Delta\mu(k,a)]\delta(k,a)$ (5.83)
@@
$u_{i}(x,a)=-aH\int\frac{d^{3}k}{(2\pi)^{3}}e^{ik\cdot x}\left[\hat{k}_{i}+\Delta\mu(k,a)T_{ij}(\hat{k})\hat{k}_{j}\right]\frac{\delta(k,a)}{k}$ (5.84)
@@
$\langle u^{2}(R)\rangle=a^{2}H^{2}\int\frac{d^{3}k}{(2\pi)^{3}}[1+\Delta\mu(k,a)]^{2}\frac{P_{\delta\delta}(k,a)}{k^{2}}W^{2}(kR)$ (5.85)
@@
$\rho_{\text{TD}}(k,a)=\Delta\mu(k,a)\bar{\rho}\delta(k,a)$ (5.86)
@@
Finally, another avenue to probe teleodynamics is by considering gravitational wave signatures. It is so as in a teleodynamic cosmology, tensor perturbations propagate not on a purely FLRW background but on an effective spacetime whose evolution includes the homogeneous component of the bias functional, as well as a possible response of the inhomogeneous sector. This implies that the propagation equation for gravitational waves will in general acquire additional friction and source terms beyond those present in general relativity and to see this, one may consider a perturbed metric of the form $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$ with $h_{\mu\nu}$ transverse and traceless. Varying the effective action which contains the term $\alpha\int d^{4}x\sqrt{-g}\,\Phi(t,x)$, leads to a modified tensor equation that may be written schematically in Fourier space as
$h^{\prime\prime}_{ij}+\left[2+\nu_{\rm TD}(a)\right]H\,h^{\prime}_{ij}+c_{T}^{2}k^{2}h_{ij}=\Pi^{\rm TD}_{ij}$ (5.87)
where we note that the primes denote derivatives with respect to conformal time, the quantity $\nu_{\rm TD}(a)$ encodes the teleodynamic correction to the effective friction term arising from the homogeneous component $\bar{\Phi}(t)$ and $\Pi^{\rm TD}_{ij}$ represents the tensorial response of the inhomogeneous part of the bias functional. Even if the propagation speed remains luminal, the presence of $\nu_{\rm TD}(a)$ modifies the evolution of the wave amplitude as it redshifts, generating a scale and time-dependent damping that depends on $d\bar{\Phi}/d\ln a$. This leads to a distinction between the electromagnetic and gravitational luminosity distances, since the strain amplitude of a gravitational wave from a standard siren at redshift $z$ will scale as
$h\propto\frac{1}{d_{L}^{\rm GW}(z)}\exp\!\left[-\frac{1}{2}\int_{0}^{z}\nu_{\rm TD}(z^{\prime})\frac{dz^{\prime}}{1+z^{\prime}}\right]$ (5.88)
This goes to imply that $d_{L}^{\rm GW}(z)$ generally differs from the electromagnetic distance $d_{L}^{\rm EM}(z)$ and the departure between these two distance measures traces the integrated history of the teleodynamic friction term and therefore serves as a direct observational probe of $\bar{\Phi}(t)$. Unlike modified gravity models with a running Planck mass, the origin of this effect lies not in a new scalar degree of freedom but in the non-equilibrium, memory-bearing nature of the cosmic medium encoded in the bias functional.
@@ 6 Law of Universal Arbitrage Equilibrium
$S_{\rm TD}(t)=\frac{C_{\rm TD}(t)}{H(t)^{2}}+\int^{t}\sigma_{\rm TD}(t^{\prime})\,dt^{\prime},\qquad T_{\rm TD}(t)=\left(\frac{\partial S}{\partial E_{H}}\right)^{-1}\simeq\frac{2\pi M_{P}^{2}}{C_{\rm TD}(t)}H(t)$ (6.1)
@@
$\dot{S}_{\rm TD}=-\frac{2C_{\rm TD}\dot{H}}{H^{3}}+\frac{\dot{C}_{\rm TD}}{H^{2}}+\sigma_{\rm TD}$ (6.2)
@@
$\dot{H}=-\frac{\rho_{\rm eff}+p_{\rm eff}}{2M_{P}^{2}}+\frac{H}{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+\frac{H^{3}}{2C_{\rm TD}}\,\sigma_{\rm TD}$ (6.5)
@@
Now, if we go by multiplying the equation (6.5) by $2H$ and using $\tfrac{d}{dt}(H^{2})=2H\dot{H}$ together with the continuity equation $\dot{\rho}_{\rm eff}+3H(\rho_{\rm eff}+p_{\rm eff})=0$ then this leads us to
$\frac{d}{dt}(H^{2})=\frac{1}{3M_{P}^{2}}\dot{\rho}_{\rm eff}+H^{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+H^{4}\frac{\sigma_{\rm TD}}{C_{\rm TD}}$ (6.6)
@@
$H^{2}(t)=\frac{\rho_{\rm eff}(t)}{3M_{P}^{2}}+\left[H^{2}(t_{\star})-\frac{\rho_{\rm eff}(t_{\star})}{3M_{P}^{2}}\right]+\int_{t_{\star}}^{t}\left(H^{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+H^{4}\frac{\sigma_{\rm TD}}{C_{\rm TD}}\right)dt^{\prime}$ (6.7)
@@
In this language, equation (6.5) becomes
$\dot{H}=-\frac{\rho_{m}+\rho_{r}+p_{r}}{2M_{P}^{2}}-\frac{\rho_{\rm TD}+p_{\rm TD}}{2M_{P}^{2}}+\frac{H}{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+\frac{H^{3}}{2C_{\rm TD}}\,\sigma_{\rm TD}$ (6.8)
revealing that two distinct teleodynamic mechanisms contribute to the expansion rate with a bulk component carried by $(\rho_{\rm TD},p_{\rm TD})=\alpha(\overline{\Phi},\overline{\Pi}_{\rm TD})$ and a horizon component encoded in $C_{\rm TD}(t)$ and $\sigma_{\rm TD}(t)$ representing the evolution of the correlation scale and the non-equilibrium memory production on the horizon. If late-time acceleration is primarily driven by bulk bias, then $C_{\rm TD}$ remains approximately constant and $\sigma_{\rm TD}$ approaches zero, with $w_{\rm TD}\approx-1$ emerging naturally when $\overline{\Pi}_{\rm TD}\simeq-\overline{\Phi}$. If instead part of the acceleration originates from the horizon sector, then the integral terms in the teleodynamic Friedmann relation are nonzero, and one infers small but finite values for $\dot{C}_{\rm TD}/C_{\rm TD}$ or $\sigma_{\rm TD}$ at low redshift. Since $\sigma_{\rm TD}\geq 0$, the last term in (6.5) always drives $H^{2}$ upward, temporarily mimicking an effective equation of state below 1-1 without introducing ghosts.
@@
$\dot{S}_{\rm TD}=-2C_{\rm TD}\frac{\dot{H}}{H^{3}}+\sigma_{\rm TD},\qquad\dot{E}_{H}=-4\pi M_{P}^{2}\frac{\dot{H}}{H^{2}},\qquad\dot{V}_{H}=-4\pi\frac{\dot{H}}{H^{4}}$ (6.10)
@@
$p_{\rm eff}=\frac{\kappa}{4\pi}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD},\qquad\kappa=\frac{2\pi M_{P}^{2}}{C_{\rm TD}}$ (6.11)
@@
$\alpha\overline{\Pi}_{\rm TD}=\frac{\kappa}{4\pi}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}=\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}$ (6.12)
@@
$\alpha\overline{\Phi}=-2M_{P}^{2}\dot{H}-(\rho_{m}+p_{m}+\rho_{r}+p_{r})-\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}$ (6.13)
@@
$p_{\rm TD}=\alpha\overline{\Pi}_{\rm TD}=-\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{|\dot{H}|}\sigma_{\rm TD}<0$ (6.14)
showing that non-equilibrium teleodynamic entropy production necessarily gives us negative pressure, and a combined expression for the EOS parameter follows
$w_{\rm TD}=\frac{p_{\rm TD}}{\rho_{\rm TD}}=\frac{-\dfrac{M_{P}^{2}}{2C_{\rm TD}}\dfrac{H^{5}}{\dot{H}}\sigma_{\rm TD}}{-2M_{P}^{2}\dot{H}-(\rho_{m}+p_{m}+\rho_{r}+p_{r})-\dfrac{M_{P}^{2}}{2C_{\rm TD}}\dfrac{H^{5}}{\dot{H}}\sigma_{\rm TD}}$ (6.15)
@@
$\sigma_{\rm TD}(t)=2\xi\frac{-\dot{H}}{H^{3}},\qquad\xi>0$ (6.16)
@@
$p_{\rm TD}=-\frac{\kappa}{2\pi}\xi H^{2}=-\frac{M_{P}^{2}}{C_{\rm TD}}\xi H^{2},\qquad\rho_{\rm TD}=\frac{3M_{P}^{2}H^{2}}{8\pi}\Big(\cdots\Big)+\frac{M_{P}^{2}}{C_{\rm TD}}\xi H^{2}$ (6.17)
@@
depends exclusively on horizon-level quantities such as $C_{\rm TD}$, $\sigma_{\rm TD}$, $H$, and $\dot{H}$ and none of the spatially varying components of the teleodynamic memory field enter this expression, indicating that cosmic acceleration arises only from the homogeneous part of the memory sector. Global memory, therefore, generates dark energy behavior, while local memory enters through spatial gradients, such as $\alpha\nabla\varphi$, and drives clustering, thereby reproducing dark matter-like dynamics.